Suppose the payment are only $16,000 each, but they are made every six months, starting six months from now. what would be the future value if the ten payments were invested at 10.0 percent annual interest? if they were invested at BankSouth at 10.0 percent compounded semiannually?
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
10 = ((1+Stated rate%/(2*100))^2-1)*100
Stated rate% = 9.76
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 16000*(((1+ 9.76/200)^(5*2)-1)/(9.76/200))
FV = 200122.67
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 16000*(((1+ 10/200)^(5*2)-1)/(10/200))
FV = 201246.28